# Formal Analysis of the Contract Automata Runtime Environment with Uppaal: Modelling, Verification and Testing

### v2

@@ 4.1. Modelling TCP/IP sockets communication
Java TCP/IP sockets communications are asynchronous with FIFO buffers [56]. In Uppaal, the interactions are via channel synchronisation and global variables. Thus, TCP/IP sockets are solely employed in the actual implementation and are not utilized by the automata of the model. Instead, in the model each party communicates with the partner using two global arrays (one for sending and one for receiving, respectively). Figure 5 shows the code used in the Uppaal models (see Section 5) for modelling the TCP/IP socket communications.
@@
The default mode for Java TCP/IP sockets is blocking [45], meaning that the sender blocks when the buffer of the receiver is full, and waits until there is enough space to proceed. Accordingly, a transition having a send in its effect will check in its guard whether there is enough space left in the array of the partner by calling either the method available (returning the space left) or isFull. Similarly, in Java TCP/IP sockets the read operation is blocking. Accordingly, before reading it is always checked whether the array is not empty with the method !isEmpty. When the array is empty the automaton blocks until a message is received.
@@ 4.2. Abstractions
As stated above, the underlying application executed by CARE is abstracted away. Therefore, properties related to the underlying application are not verifiable under the current abstraction. An example of property not verifiable under the current abstraction could be to show that the returned payload of an action (e.g., an Integer object) is always non-negative, by reasoning on the code of the corresponding method returning the value. This kind of properties cannot be verified under the current abstraction. In this case, a theorem prover like KeY [2] would be more suited than Uppaal.
@@ 4.3. Traceability
This paper adopts a lightweight approach to integrate the formal model with the concrete implementation, using the facilities provided by Uppaal. Traceability and model-based testing have been used to ascertain the adequacy of the model with respect to the implementation. Traceability involves connecting the items of the abstract model with the lines of source code that are being abstractly represented. As stated previously, some aspects of the application are abstracted away. Therefore, only the relevant source code lines are connected with the abstract model. In [10] each transition of the model is traced back to the corresponding source code instructions using comments in the model. Traceability information is related to a specific version of the source code [^1] https://github.com/contractautomataproject/CARE/releases/tag/v1.0.1. For example, Figure 6 shows a portion of the automaton modelling the RunnableOrchestratedContract (discussed in details in the next section). Specifically, from state Ready a transition is executed which is traced back to line 98 of the class RunnableOrchestratedContract, also showed in Figure 6. Indeed, at line 98 the first read operation from the ObjectInputStream, reading from the TCP/IP socket, is performed by the service. This is abstracted in the model by the operation dequeue() of the automaton.
@@ 5. Formal model
This section describes the formal model of CARE. All models used in this paper together with the evaluated formulas are available in [10]. The network of automata is composed of a RunnableOrchestration automaton modelling the orchestrator, and each service is modeled by two automata: RunnableOrchestratedContract and SocketTimeout. Figure 7 shows the system declarations of the model. In particular, there is a single instance of RunnableOrchestration whilst there are N instance of RunnableOrchestratedContract and SocketTimeout. Indeed, both RunnableOrchestratedContract and SocketTimeout templates take as parameter id_t, which is a special global declaration of Uppaal (namely, typedef int[0,N-1] id_t) to instruct the model to produce N replicas of these template, where as stated above N is also a global constant of the model.
@@
Figure 9 displays the template automaton for the RunnableOrchestration (i.e., the orchestrator), while the template automaton for the RunnableOrchestratedContract (i.e., the service) has as parameter the id of the service (of type id_t) and is depicted in Figure 10. Recall that, for readability, the Timeout location is not displayed (see Section 4.1). The behaviour according to the given configuration of action and choice is modelled inside each automaton. We anticipate that we will use variants of these two automata in Section 6, in order to perform different analyses. In particular, in Figure 9 the configurations of choice and action (for both the services and the orchestrator) are instantiated non-deterministically by the transition outgoing the initial state (see Section 5.1). Note that, in this case, the variables choice and action are not parameters of the templates. Another version of the model is also available, where the configuration options, instead of being selected non-deterministically, are fixed as parameters of the templates. Note that in Uppaal templates cannot be primitively instantiated non-deterministically. The repository [10] contains the model equipped with traceability information, i.e., each transition of the model has a comment describing the class and the lines of the source code that correspond to the specific behaviour of the transition. The models used for generating tests, together with the generated tests, are also available in [10].
@@ 6.1. Validation through modelling
The first validation was performed during the modelling phase. Indeed, formal modelling requires an accurate analysis and review of the source code. Interactive simulation is used during modelling to animate and analyse the portion of the model designed so far, similarly to how the source code is debugged interactively (e.g., by choosing the next step). We note that in many model-based engineering tools, behavioural models (e.g., state charts) are validated by only relying on graphical interactive simulations [22]. Issues can be detected during this phase in particular if the source code has not been thoroughly tested, as was the case for CARE.
@@
On a side note, thorough testing is generally more time-consuming than designing a formal model similar to the one in this paper. This is a further benefit derived from the usage of formal methods. For example, the library implementing contract automata operations (CATLib[13] has been tested up to 100% coverage of all lines and branches, also using mutation testing [13]. In CATLib, the lines of code of the tests are more than three times those of its source code [13]. A similar effort for CARE is more demanding than the one needed for designing the models in Figures 9 and 10. Uppaal has been used to automatically generate tests from the formal model.
@@ 6.2. Formal Verification
Performances. Statistical model checking has been used to scale to larger systems, and the verification has been performed in a few seconds on a standard laptop. Conversely, exhaustive model checking necessitated more resources and has been limited to smaller parameter setup generating hundreds of millions of states (see below). In this case, the verification has been performed on a machine with Intel(R) Core(TM) i9-9900K CPU @ 3.60 GHz equipped with 32 GB of RAM. Uppaal has been configured for maximising the state space optimisation and reusing the generated state space. Logs of the experiments are available in [10]. The summaries of the analysis performed with statistical (resp., exhaustive) model checking are in Table 1 (resp., Table 2).
@@ 6.3. Parameters Tuning
In this section, when reporting the evaluation of a probabilistic formula, instead of reporting an interval containing zero (e.g., $[0,0.00996915]$), sometimes, for conciseness, we will say that the property has a probability near zero. To be more precise, we mean that all runs do not satisfy the property, thus the probability lies within the interval $[0,0+\varepsilon]$ with probability $1-\alpha$. We remind readers that, unlike probabilistic model checking, in statistical model checking, the unknown probability value is estimated to lie within an interval with a given confidence.
@@
Figure 11. On the left side, the SocketTimeout template automaton (one such automaton is istantiated for each service). On the right side, the orchestration automaton taken from the composition service example in [15]
@@
which measures the probability that, within 500 time units, one of the arrays orc2services[i] becomes full (ror is the orchestrator automaton). In three separate experiments where queueSize was varied between 3, 4, and 5, this formula yielded the respective evaluation intervals of $[0.990031,1]$, $[0.00632077,0.0163207]$ (with $\alpha\!=\!0.005$), and $[0,0.00996915]$. Based on these results, if not stated otherwise, queueSize is set to 5 for the subsequent experiments.
@@
(ROC is used as an abbreviation of RunnableOrchestratedContract). In different experiments where we varied the values of the pair (rate,timeout), specifically in $(5,14)$, $(4,15)$, $(5,15)$, respectively, the first formula (timeout probability) yielded the respective evaluation intervals $[3.06006e^{-06},0.00999494]$ (with $\alpha\!=\!0.005$), $[0.0433422,0.053341]$, and $[0,0.00996915]$, while the second formula (probability of termination) yielded $[0.990005,0.999997]$ (with $\alpha\!=\!0.005$), $[0.948625,0.958625]$ (with $\alpha\!=\!0.005$), and $[0.990031,1]$. Therefore, to fulfill the aforementioned three objectives, we have set the values of write and read to 5, while the value of timeout has been set to 15 for the subsequent experiments. Indeed, in the experiments we just mentioned, the value (5,15) for the pair (rate,timeout) yields the best values for both the probability of experiencing timeout and the probability of reaching termination.
@@ 6.4. Verification
ror.Stop-->((ror.Terminated&&(forall(i:id_t)ROC(i).Terminated))||
@@
Note that this kind of problems are hard to detect without model checking. Indeed, the counterexample trace was generated automatically, and the counterexample trace is composed of hundreds of steps. Without model checking this would require to manually execute each step of this trace, and the longer the trace the less chance to discover it. After we detected this issue, we analysed the underlying Java TCP/IP socket semantics [45] and fixed the model as described in Section 4 (i.e., by modelling these sockets with default blocking mode). Another fix could be to include an ack after the reception of a SKIP message (this would however require to modify also the implementation).
@@
We also include in the repository [10] a version of the model where each pair of services has its own buffers. All results in this section also hold in that model.
@@ 7.3. An example of test generation
All versions of the model contains queries for testing the various configurations of the orchestrator (e.g., majoritarian choice, distributed action). We refer to [10] for more details about these other queries. In the concrete tests, the orchestrator will interact with two mocked services. Furthermore, the orchestration contract automaton, a parameter of the RunnableOrchestration object (see Figure 3), is dummy and is instrumental to cover the desired portions of code of CARE. In particular, the orchestration contract automaton that will be used in the concrete tests accepts the trace (!euro,-)(?coffee,!coffee)[(!euro,-)]. The corresponding query used to generate the abstract test is:
@@ 7.3.1. Test generation
{minted}
[mathescape,fontsize=,linenos]java msg = (String) oin.readObject();
//INIT CHECK assertEquals(msg,RunnableOrchestration.check_msg);
String received_choiceType = (String) oin.readObject(); String received_actionType = (String) oin.readObject(); assertEquals(received_actionType,actionType); assertEquals(received_choiceType,choiceType);
oout.writeObject(RunnableOrchestration.ack_msg); oout.flush(); //STOP CHECK
@@
We now move to the structure of the concrete JUnit test and the concretisation of the abstract test code discussed until now. After that, we will discuss the remaining part of the abstract test code and its concretisation. The concrete test file is available at [10] and is called DictatorialCentralisedRunnableOrchestrationTest.java. All concrete tests used for testing the orchestrator follow the same schema, reported below (for the case of dictatorial choice and centralised action).
{listing}
@@
{minted}
[mathescape,fontsize=,linenos]java @Test public void test() throws IOException, ClassNotFoundException, InterruptedException Thread alice = new Thread(()-¿ //mocked service check(8080); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8080)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_alice(oin,oout);
catch (IOException — ClassNotFoundException e) throw new RuntimeException(e); );
Thread bob = new Thread(()-¿ //mocked service check(8081); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8081)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_bob(oin,oout); catch (IOException — ClassNotFoundException e) throw new RuntimeException(e); ); alice.start(); bob.start(); RunnableOrchestration ror = new DictatorialChoiceRunnableOrchestration(new Agreement(), adc.importMSCA(dir+”Orchestration.data”), Arrays.asList(null,null), Arrays.asList(8080,8081), new CentralisedOrchestratorAction());
Thread tror = new Thread(ror); tror.start(); tror.join();
@@
{minted}
[mathescape,fontsize=,linenos]java private void uppaal_check( ObjectInputStream oin, ObjectOutputStream oout) throws IOException String choiceType = ”Dictatorial”; String actionType = ”Centralised”; String msg; msg = (String) oin.readObject(); assertEquals(msg,RunnableOrchestration.check_msg); String received_choiceType = (String) oin.readObject(); String received_actionType = (String) oin.readObject(); assertEquals(received_actionType,actionType); assertEquals(received_choiceType,choiceType); oout.writeObject(RunnableOrchestration.ack_msg); oout.flush();
@@
{minted}
[mathescape,fontsize=,linenos]java //centralised offerer 0
msg = (String) oin.readObject();
assertEquals(msg, expected action);
msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,expectedPayload);
oout.writeObject(INSERT OFFER); oout.flush();
The transition outgoing state CentralisedOffer of the RunnableOrchestration automaton (see Figure 9) has as test code //centralised offerer \$(ror.offerer). This will append the comment at line 1 in the abstract test, indicating that the next instructions of the abstract test will cover the first offer action performed by Alice (service at index 00). This is helpful for generating the concrete test from the abstract test. After performing the check, both Alice and Bob are again in state Ready (Figure 12). The first readObject operation (line 3) is again appended by the outgoing transition from state Ready of Alice. The next transition executed by Alice, with guard d1==ACTION, will append the instruction at line 5. The assertion will check whether the message received by the orchestrator will be the expected action, which will be concretised in the concrete test with the value "euro" (i.e., the orchestrator is invoking the method euro of Alice).
@@
{minted}
[mathescape,fontsize=,linenos]java //requester: 0
msg = (String) oin.readObject();
assertEquals(msg, expected action);
msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,expectedPayload);
oout.writeObject(INSERT REQUEST); oout.flush();
//matching offerer: 1
msg = (String) oin.readObject();
assertEquals(msg, expected action);
msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,expectedPayload);
oout.writeObject(INSERT OFFER); oout.flush();
@@
msg = (String) oin.readObject(); //receiving the matching offer assertEquals(msg, expected payload);
@@
{minted}
[mathescape,fontsize=,linenos]java msg = (String) oin.readObject();
assertEquals(msg,RunnableOrchestration.choice_msg);
@@
{minted}
[mathescape,fontsize=,linenos]java msg = (String) oin.readObject();
assertEquals(msg,RunnableOrchestration.stop_msg);
@@
{minted}
[mathescape,fontsize=,linenos]java private void uppaal_alice( ObjectInputStream oin, ObjectOutputStream oout) throws IOException String msg; msg = (String) oin.readObject(); assertEquals(msg, ”euro”); msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,null); oout.writeObject(null); oout.flush();
msg = (String) oin.readObject(); assertEquals(msg, ”coffee”); msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,null); oout.writeObject(”request payload”); oout.flush();
msg = (String) oin.readObject(); //receiving the matching offer assertEquals(msg, ”offer payload”);
msg = (String) oin.readObject(); assertEquals(msg,RunnableOrchestration.choice_msg);
@@
//if the orchestrator decides not to terminate while (msg.equals(”euro”)) msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg, null); oout.writeObject(null); oout.flush(); msg = (String) oin.readObject(); assertEquals(msg, RunnableOrchestration.choice_msg); msg = (String) oin.readObject();
assertEquals(msg,RunnableOrchestration.stop_msg);
@@
{minted}
[mathescape,fontsize=,linenos]java private void uppaal_bob(ObjectInputStream oin, ObjectOutputStream oout) throws IOException String msg;
msg = (String) oin.readObject(); assertEquals(msg, ”coffee”); msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,”request payload”); oout.writeObject(”offer payload”); oout.flush();
msg = (String) oin.readObject(); assertEquals(msg,RunnableOrchestration.choice_msg);
//in case the orchestrator decides not to terminate do msg = (String) oin.readObject(); while (msg.equals(RunnableOrchestration.choice_msg));
assertEquals(msg,RunnableOrchestration.stop_msg);
@@
This concludes the example of test generation. Recall that this is one query of one version of the model (for testing the orchestrator). There are other queries used for testing the other configurations (e.g., majoritarian choice, distributed offer) and other versions of the model for testing the other components. However, the generation of abstract tests and their concretisation follows the same approach covered by this example. All these versions and the concrete tests are available at [10].

### v3

@@ 4.1. Modelling TCP/IP sockets communication
Java TCP/IP sockets communications are asynchronous with FIFO buffers [Obj]. In Uppaal, the interactions are via channel synchronisation and global variables. Thus, TCP/IP sockets are solely employed in the actual implementation and are not utilized by the automata of the model. Instead, in the model each party communicates with the partner using two global arrays (one for sending and one for receiving, respectively). Figure 5 shows the code used in the Uppaal models (see Section 5) for modelling the TCP/IP socket communications.
@@
The default mode for Java TCP/IP sockets is blocking [Jav], meaning that the sender blocks when the buffer of the receiver is full, and waits until there is enough space to proceed. Accordingly, a transition having a send in its effect will check in its guard whether there is enough space left in the array of the partner by calling either the method available (returning the space left) or isFull. Similarly, in Java TCP/IP sockets the read operation is blocking. Accordingly, before reading it is always checked whether the array is not empty with the method !isEmpty. When the array is empty the automaton blocks until a message is received.
@@ 4.2. Abstractions
As stated above, the underlying application executed by CARE is abstracted away. Therefore, properties related to the underlying application are not verifiable under the current abstraction. An example of property not verifiable under the current abstraction could be to show that the returned payload of an action (e.g., an Integer object) is always non-negative, by reasoning on the code of the corresponding method returning the value. This kind of properties cannot be verified under the current abstraction. In this case, a theorem prover like KeY [ABB+16] would be more suited than Uppaal.
@@ 4.3. Traceability
This paper adopts a lightweight approach to integrate the formal model with the concrete implementation, using the facilities provided by Uppaal. Traceability and model-based testing have been used to ascertain the adequacy of the model with respect to the implementation. Traceability involves connecting the items of the abstract model with the lines of source code that are being abstractly represented. As stated previously, some aspects of the application are abstracted away. Therefore, only the relevant source code lines are connected with the abstract model. In [Basb] each transition of the model is traced back to the corresponding source code instructions using comments in the model. Traceability information is related to a specific version of the source code [^1] https://github.com/contractautomataproject/CARE/releases/tag/v1.0.1. For example, Figure 6 shows a portion of the automaton modelling the RunnableOrchestratedContract (discussed in details in the next section). Specifically, from state Ready a transition is executed which is traced back to line 98 of the class RunnableOrchestratedContract, also showed in Figure 6. Indeed, at line 98 the first read operation from the ObjectInputStream, reading from the TCP/IP socket, is performed by the service. This is abstracted in the model by the operation dequeue() of the automaton.
@@ 5. Formal model
This section describes the formal model of CARE. All models used in this paper together with the evaluated formulas are available in [Basb]. The network of automata is composed of a RunnableOrchestration automaton modelling the orchestrator, and each service is modeled by two automata: RunnableOrchestratedContract and SocketTimeout. Figure 7 shows the system declarations of the model. In particular, there is a single instance of RunnableOrchestration whilst there are N instance of RunnableOrchestratedContract and SocketTimeout. Indeed, both RunnableOrchestratedContract and SocketTimeout templates take as parameter id_t, which is a special global declaration of Uppaal (namely, typedef int[0,N-1] id_t) to instruct the model to produce N replicas of these template, where as stated above N is also a global constant of the model.
@@
Figure 9 displays the template automaton for the RunnableOrchestration (i.e., the orchestrator), while the template automaton for the RunnableOrchestratedContract (i.e., the service) has as parameter the id of the service (of type id_t) and is depicted in Figure 10. Recall that, for readability, the Timeout location is not displayed (see Section 4.1). The behaviour according to the given configuration of action and choice is modelled inside each automaton. We anticipate that we will use variants of these two automata in Section 6, in order to perform different analyses. In particular, in Figure 9 the configurations of choice and action (for both the services and the orchestrator) are instantiated non-deterministically by the transition outgoing the initial state (see Section 5.1). Note that, in this case, the variables choice and action are not parameters of the templates. Another version of the model is also available, where the configuration options, instead of being selected non-deterministically, are fixed as parameters of the templates. Note that in Uppaal templates cannot be primitively instantiated non-deterministically. The repository [Basb] contains the model equipped with traceability information, i.e., each transition of the model has a comment describing the class and the lines of the source code that correspond to the specific behaviour of the transition. The models used for generating tests, together with the generated tests, are also available in [Basb].
@@ 6.1. Validation through modelling
The first validation was performed during the modelling phase. Indeed, formal modelling requires an accurate analysis and review of the source code. Interactive simulation is used during modelling to animate and analyse the portion of the model designed so far, similarly to how the source code is debugged interactively (e.g., by choosing the next step). We note that in many model-based engineering tools, behavioural models (e.g., state charts) are validated by only relying on graphical interactive simulations [BMF23]. Issues can be detected during this phase in particular if the source code has not been thoroughly tested, as was the case for CARE.
@@
On a side note, thorough testing is generally more time-consuming than designing a formal model similar to the one in this paper. This is a further benefit derived from the usage of formal methods. For example, the library implementing contract automata operations (CATLib[BtB22] has been tested up to 100% coverage of all lines and branches, also using mutation testing [BtB22]. In CATLib, the lines of code of the tests are more than three times those of its source code [BtB22]. A similar effort for CARE is more demanding than the one needed for designing the models in Figures 9 and 10. Uppaal has been used to automatically generate tests from the formal model.
@@ 6.2. Formal Verification
Performances. Statistical model checking has been used to scale to larger systems, and the verification has been performed in a few seconds on a standard laptop. Conversely, exhaustive model checking necessitated more resources and has been limited to smaller parameter setup generating hundreds of millions of states (see below). In this case, the verification has been performed on a machine with Intel(R) Core(TM) i9-9900K CPU @ 3.60 GHz equipped with 32 GB of RAM. Uppaal has been configured for maximising the state space optimisation and reusing the generated state space. Logs of the experiments are available in [Basb]. The summaries of the analysis performed with statistical (resp., exhaustive) model checking are in Table 1 (resp., Table 2).
@@ 6.3. Parameters Tuning
In this section, when reporting the evaluation of a probabilistic formula, instead of reporting an interval containing zero (e.g., $[0,0.00996915]$), sometimes, for conciseness, we will say that the property has a probability near zero. To be more precise, we mean that all runs do not satisfy the property, thus the probability lies within the interval $0,0+\varepsilon]$ with probability $1-\alpha$. We remind readers that, unlike probabilistic model checking, in statistical model checking, the unknown probability value is estimated to lie within an interval with a given confidence.
@@
Figure 11. On the left side, the SocketTimeout template automaton (one such automaton is istantiated for each service). On the right side, the orchestration automaton taken from the composition service example in [BtB23a]
@@
which measures the probability that, within 500 time units, one of the orc2services[i] arrays becomes full (ror is the orchestrator automaton). In three separate experiments where queueSize was varied between 3, 4, and 5, this formula yielded the respective evaluation intervals of $[0.990031,1]$, $[0.00632077,0.0163207]$ (with $\alpha\!=\!0.005$), and $[0,0.00996915]$. Based on these results, if not stated otherwise, queueSize is set to 5 for the subsequent experiments.
@@
(ROC is used as an abbreviation of RunnableOrchestratedContract). In different experiments where we varied the values of the pair (rate,timeout), specifically in $(5,14)$, $(4,15)$, $(5,15)$, respectively, the first formula (timeout probability) yielded the respective evaluation intervals $3.06006e^{-06},0.00999494]$ (with $\alpha\!=\!0.005$), $[0.0433422,0.053341]$, and $[0,0.00996915]$, while the second formula (probability of termination) yielded $[0.990005,0.999997]$ (with $\alpha\!=\!0.005$), $[0.948625,0.958625]$ (with $\alpha\!=\!0.005$), and $[0.990031,1]$. Therefore, to fulfill the aforementioned three objectives, we have set the values of write and read to 5, while the value of timeout has been set to 15 for the subsequent experiments. Indeed, in the experiments we just mentioned, the value (5,15) for the pair (rate,timeout) yields the best values for both the probability of experiencing timeout and the probability of reaching termination.
@@ 6.4. Verification
ror.Stop-->((ror.Terminated&&(forall(i:id_t)ROC(i).Terminated))$|$
@@
Note that this kind of problems are hard to detect without model checking. Indeed, the counterexample trace was generated automatically, and the counterexample trace is composed of hundreds of steps. Without model checking this would require to manually execute each step of this trace, and the longer the trace the less chance to discover it. After we detected this issue, we analysed the underlying Java TCP/IP socket semantics [Jav] and fixed the model as described in Section 4 (i.e., by modelling these sockets with default blocking mode). Another fix could be to include an ack after the reception of a SKIP message (this would however require to modify also the implementation).
@@
We also include in the repository [Basb] a version of the model where each pair of services has its own buffers. All results in this section also hold in that model.
@@ 7.3. An example of test generation
All versions of the model contains queries for testing the various configurations of the orchestrator (e.g., majoritarian choice, distributed action). We refer to [Basb] for more details about these other queries. In the concrete tests, the orchestrator will interact with two mocked services. Furthermore, the orchestration contract automaton, a parameter of the RunnableOrchestration object (see Figure 3), is dummy and is instrumental to cover the desired portions of code of CARE. In particular, the orchestration contract automaton that will be used in the concrete tests accepts the trace (!euro,-)(?coffee,!coffee)[(!euro,-)]. The corresponding query used to generate the abstract test is:
@@ 7.3.1. Test generation
msg = (String) oin.readObject();
[…]
assertEquals(msg,RunnableOrchestration.check_msg);
[…]
String received_actionType = (String) oin.readObject();
assertEquals(received_actionType,actionType);
assertEquals(received_choiceType,choiceType);
[…]
oout.flush();
//STOP CHECK
@@
We now move to the structure of the concrete JUnit test and the concretisation of the abstract test code discussed until now. After that, we will discuss the remaining part of the abstract test code and its concretisation. The concrete test file is available at [Basb] and is called DictatorialCentralisedRunnableOrchestrationTest.java. All concrete tests used for testing the orchestrator follow the same schema, reported below (for the case of dictatorial choice and centralised action).
@@
@Test
public void test() throws IOException, ClassNotFoundException, InterruptedException {
Thread alice = new Thread(()->{ //mocked service
check($80$);
try (ServerSocketChannel server = ServerSocketChannel.open()){
server.bind(new InetSocketAddress($80$));
Socket socket = server.accept().socket();
ObjectInputStream oin = new ObjectInputStream(socket.getInputStream());
ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream());
oout.flush();
uppaal_alice(oin,oout);
[…]
throw new RuntimeException(e);
[…]
check(8081);
try (ServerSocketChannel server = ServerSocketChannel.open()){
server.bind(new InetSocketAddress(8081));
Socket socket = server.accept().socket();
ObjectInputStream oin = new ObjectInputStream(socket.getInputStream());
ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream());
oout.flush();
uppaal_bob(oin,oout);
} catch (IOException | ClassNotFoundException e) {
throw new RuntimeException(e);
alice.start();
bob.start();
RunnableOrchestration ror = new DictatorialChoiceRunnableOrchestration(new Agreement(),
adc.importMSCA(dir+”Orchestration.data”),
Arrays.asList(null,null),
Arrays.asList(8080,8081),
new CentralisedOrchestratorAction());
[…]
tror.start();
tror.join();
@@
private void uppaal_check( ObjectInputStream oin, ObjectOutputStream oout) throws IOException {
String choiceType = Dictatorial”;
String actionType = Centralised”;
String msg;
msg = (String) oin.readObject();
assertEquals(msg,RunnableOrchestration.check_msg);
String received_choiceType = (String) oin.readObject();
String received_actionType = (String) oin.readObject();
assertEquals(received_actionType,actionType);
assertEquals(received_choiceType,choiceType);
oout.writeObject(RunnableOrchestration.ack_msg);
oout.flush();
@@
//centralised offerer 0
[…]
assertEquals(msg,expectedPayload);
[…]
oout.flush();
The transition outgoing state CentralisedOffer of the RunnableOrchestration automaton (see Figure 9) has as test code //centralised offerer \$(ror.offerer). This will append the comment at line 1 in the abstract test, indicating that the next instructions of the abstract test will cover the first offer action performed by Alice (service at index 0). This is helpful for generating the concrete test from the abstract test. After performing the check, both Alice and Bob are again in state Ready (Figure 12). The first readObject operation (line 3) is again appended by the outgoing transition from state Ready of Alice. The next transition executed by Alice, with guard d1==ACTION, will append the instruction at line 5. The assertion will check whether the message received by the orchestrator will be the expected action, which will be concretised in the concrete test with the value "euro" (i.e., the orchestrator is invoking the method euro of Alice).
@@
[…]
assertEquals(msg,expectedPayload);
[…]
oout.flush();
[…]
assertEquals(msg,expectedPayload);
[…]
oout.flush();
[…]
assertEquals(msg, expected payload);
@@
msg = (String) oin.readObject();
[…]
@@
msg = (String) oin.readObject();
[…]
@@
private void uppaal_alice( ObjectInputStream oin, ObjectOutputStream oout) throws IOException {
String msg;
@@
assertEquals(msg, euro”);
msg = (String) oin.readObject(); // reading payload centralised action
assertEquals(msg,null);
oout.writeObject(null);
oout.flush();
[…]
assertEquals(msg, coffee”);
msg = (String) oin.readObject(); // reading payload centralised action
assertEquals(msg,null);
oout.writeObject(”request payload”);
oout.flush();
[…]
assertEquals(msg, offer payload”);
[…]
assertEquals(msg,RunnableOrchestration.choice_msg);
[…]
while (msg.equals(”euro”)) {
msg = (String) oin.readObject(); // reading payload centralised action
assertEquals(msg, null);
oout.writeObject(null);
oout.flush();
msg = (String) oin.readObject();
assertEquals(msg, RunnableOrchestration.choice_msg);
msg = (String) oin.readObject();
[…]
@@
private void uppaal_bob(ObjectInputStream oin, ObjectOutputStream oout) throws IOException {
String msg;
[…]
assertEquals(msg, coffee”);
msg = (String) oin.readObject(); // reading payload centralised action
assertEquals(msg,”request payload”);
oout.writeObject(”offer payload”);
oout.flush();
[…]
assertEquals(msg,RunnableOrchestration.choice_msg);
[…]
do {
msg = (String) oin.readObject();
} while (msg.equals(RunnableOrchestration.choice_msg));
[…]
@@
This concludes the example of test generation. Recall that this is one query of one version of the model (for testing the orchestrator). There are other queries used for testing the other configurations (e.g., majoritarian choice, distributed offer) and other versions of the model for testing the other components. However, the generation of abstract tests and their concretisation follows the same approach covered by this example. All these versions and the concrete tests are available at [Basb].